OpenStudy (anonymous):

Question 5 of 10 Multiple Choice: Please select the best answer and click "submit." Solve the equation 3x - 4y = 16 for x. A. x = 4y + 16 B. C. D. x = 3(16 + 4y)

OpenStudy (anonymous):

B. X= 4y/3 + 16/3 C. X= 16/3 - 4y/3

TheSmartOne (thesmartone):

Lets add 4y on both sides 3x=4y+16 Now divide by 3 x=(4/3)y+(16/3)

OpenStudy (anonymous):

Thanks you thesmartone